# Compatible Poisson Structures On Multiplicative Quiver Varieties

### v2

@@ 4. Noncommutative origin of the pencil
As a consequence of Theorem 4.5, we can repeat Remark 4.2 and see the family of Poisson brackets $(z_{0}\{-,-\}_{\mathtt{VdB}}+\{-,-\}_{\underline{z}})$ on $\operatorname{Sym}(H_{0}(A_{Q}))$ as the universal Poisson pencil that induces the one on the varieties $M_{\overline{Q}}^{\gamma}/\!/\operatorname{GL}_{\mathbf{n}}$ defined by taking invariants in Theorem 3.14. In fact, any computation can be performed using the operations ({{,}}$(\left\{\!\!\left\{-,-\right\}\!\!\right\}_{o},\left\{\!\!\left\{-,-\right\}\!\!\right\}_{r})$ on $A_{Q}$.
@@
An analogue of Theorem 4.5 in the case of character varieties (cf. § 3.3.2) can be obtained by replacing Van den Bergh’s outer double (quasi-Poisson) bracket on $A_{Q}$ by the one introduced by Massuyeau-Turaev in [31] on $\mathbb{C}\langle a_{1}^{\pm 1},(a_{1}^{\ast})^{\pm 1},\ldots,a_{g}^{\pm 1},(a_{g}^{\ast})^{\pm 1},z_{1}^{\pm 1},\ldots,z_{r}^{\pm 1}\rangle$. The case of quiver varieties (cf. § 3.3.4) is obtained similarly by using Van den Bergh’s outer double Poisson bracket on $\mathbb{C}\overline{Q}$ given in [39, §6.3-6.4]. Alternatively note that the outer double bracket on $\mathbb{C}\overline{Q}$ in the Poisson case reduces to the first constant term appearing in (4.2), thus we can formally derive this case from Theorem 4.5 by setting $\gamma_{a}=\tilde{\gamma}\in\mathbb{R}^{\times}$ for each $a\in\overline{Q}$ and considering the family $(z_{0}\tilde{\gamma}^{-1}\{-,-\}_{\mathtt{VdB}}+\{-,-\}_{\underline{z}})$ in the limit $\tilde{\gamma}\to\infty$. For generalized MQV (cf. § 3.3.3), it is conjectured in [19] that their Poisson structure is induced by an outer double bracket; as this problem is still open, exhibiting a universal Poisson pencil is currently beyond reach.
@@ 5.1. General construction
[…]
@@
[…]
@@ 5.1.1. The pencil
Due to Theorem 3.14, the bivector $P+\psi_{\underline{z}}$ is always non-degenerate quasi-Poisson on $M_{\overline{Q}_{d}}^{\gamma}$ for the moment map (5.3) (or equivalently (5.7)). In particular, this yields a quasi-Poisson pencil centered at $P$ of order $d(d-1)/2$. After performing reduction, we obtain a Poisson pencil on the spin RS phase space $\mathcal{C}_{n,d,q}^{\times}$ (5.8).
@@ 5.2. Local parametrization
We continue with the notations of the previous subsection. Following [12, §4.1], consider the affine space (×$(\mathbb{C}^{\times})^{n}\times\mathbb{C}^{2nd}$ with coordinates $(x_{i},a_{i}^{\alpha},b_{i}^{\alpha})_{1\leq i\leq n}^{1\leq\alpha\leq d}$, and define the following subspace of dimension $2nd$:
@@
is well-defined and injective. Furthermore, $\xi$ is a morphism of Poisson varieties if $\mathcal{C}_{n,d,q}^{\times}$ is endowed with the Poisson bracket defined by the Poisson bivector $P$ (3.16) and $\mathfrak{h}_{\mathrm{reg}}/S_{n}$ is endowed with the Poisson bracket {,}$\{-,-\}_{0}$ uniquely determined by
@@
[…]
@@ 5.2.1. Local expression of $\psi_{\underline{z}}$
By reduction of the quasi-Poisson pencil given in § 5.1.1, any bivector $z_{0}P+\psi_{\underline{z}}$ on $\mathcal{C}_{n,d,q}^{\times}$ with $z_{0}\in\mathbb{C}$ is Poisson, and it is non-degenerate whenever $z_{0}\neq 0$, cf. Corollary 3.10. Restricting the Poisson pencil to $\mathcal{C}^{\circ}$, the isomorphism $\xi:\mathfrak{h}_{\mathrm{reg}}/S_{n}\to\mathcal{C}^{\circ}$ that intertwines the Poisson structures {,}$\{-,-\}_{0}$ with $P$ (by Proposition 5.1) and $\psi_{\underline{z}}^{\mathrm{loc}}$ with $\psi_{\underline{z}}$ (by Lemma 5.2) is therefore intertwining the structure $z_{0}\{-,-\}_{0}+\psi_{\underline{z}}^{\mathrm{loc}}$ with $z_{0}P+\psi_{\underline{z}}$, making the former a Poisson bracket for any possible parameter. These observations can be summarized in the following form.
@@
There is a Poisson pencil on $\mathfrak{h}_{\mathrm{reg}}/S_{n}$ spanned by the Poisson brackets {,}$\{-,-\}_{0}$ and $\psi_{\underline{z}}^{\mathrm{loc}}$, with $\underline{z}=(z_{\alpha\beta})_{1\leq\alpha<\beta\leq d}$ defined as in § 5.1.1. Any linear combination $z_{0}\{-,-\}_{0}+\psi_{\underline{z}}^{\mathrm{loc}}$ with $z_{0}\in\mathbb{C}$ is therefore a Poisson bracket, which is non-degenerate when $z_{0}\neq 0$. This pencil (and therefore the pencil spanned by $P$ and $\psi_{\underline{z}}$ on $\mathcal{C}_{n,d,q}^{\times}$) has order $r\geq\frac{(d-1)(d-2)}{2}+1$.
@@
We only need to compute the bound on the order of the pencil. Since each $\psi_{\underline{z}}^{\mathrm{loc}}$ is degenerate while {,}$\{-,-\}_{0}$ is not, the order equals $\mathrm{ord}(\psi_{\underline{z}}^{\mathrm{loc}})+1$.
@@
Remark 5.5.
We expect that the inequality for the order given in Corollary 5.4 is an equality. This is the case for $d=2$, as the order is precisely $1$. To see this, note that $\psi_{\underline{z}}^{\mathrm{loc}}\equiv 0$ identically for any $z_{12}\in\mathbb{C}$ because $G_{\underline{z}}(i,j;\alpha,\beta)$ (5.17) is always vanishing on $\mathfrak{h}_{\mathrm{reg}}$.
@@ 5.2.2. Comparison with the structure of Arutyunov-Olivucci
±$\pm\frac{1}{2}\sum_{\gamma=1}^{d}o(\alpha,\gamma)b_{j}^{\beta}a_{j}^{\alpha}a_{i}^{\gamma}\pm\frac{1}{2}\left(\sum_{\gamma,\epsilon}o(\epsilon,\gamma)a_{i}^{\epsilon}a_{j}^{\gamma}\right)a_{i}^{\alpha}b_{j}^{\beta}+\frac{1}{2}a_{i}^{\alpha}a_{i}^{\beta}b_{j}^{\beta}-\frac{1}{2}\delta_{\alpha\beta}a_{i}^{\alpha}b_{j}^{\beta}\,,$ (5.24c)
@@
The minus Poisson bracket {$\{-,-\}_{-}$ constructed by Arutyunov and Olivucci [5] and the Poisson bracket {,}$\{-,-\}_{0}$ constructed by Chalykh and Fairon [12] belong to the pencil of compatible Poisson brackets from Corollary 5.4.
@@
By definition, {,}$\{-,-\}_{0}$ is certainly in the pencil. Take the sequence of parameters $\underline{z}^{-}$ given by $z_{\alpha\beta}^{-}=\frac{1}{2}$, $\alpha<\beta$. The corresponding skewsymmetric matrix has entries $\frac{1}{2}o(\alpha,\beta)$, and the Poisson structure $\psi_{\underline{z}^{-}}^{\mathrm{loc}}$ (5.2) is given in terms of
@@
It is then an exercise to check that {,}={,}$\{-,-\}_{-}=\{-,-\}_{0}+\psi_{\underline{z}^{-}}^{\mathrm{loc}}$ using (5.1) and (§ 5.2.2), remembering that $\sum_{\gamma}a_{j}^{\gamma}=1$ for each $1\leq j\leq n$. ∎
@@
Finally, we can use Proposition 5.6 to provide another quiver interpretation of the minus Poisson structure of Arutyunov and Olivucci [5]. To do so, we need to unwind the construction of the master phase space $M_{\overline{Q}_{d}}^{\gamma}$ considered in § 5.1. At a purely geometric level, the quasi-Poisson structure is obtained by considering the internally fused double $\operatorname{GL}_{n}\times\operatorname{GL}_{n}=\{(X,Z)\}$ as a $\operatorname{GL}_{n}$-variety, with $d$ copies of the space $(T^{\ast}\mathbb{C}^{n})^{\circ}:=\{(W,V)\in T^{\ast}\mathbb{C}^{n}\mid 1+VW\neq 0\}$ seen as a $(\operatorname{GL}_{n}\times\mathbb{C}^{\times})$-variety, and we perform fusion of these different actions in a specific order. Observe from Remark 3.7 that the fusion terms added by the $\mathbb{C}^{\times}$ actions will, in fact, appear as bivectors $\psi_{\underline{z}}$ in the pencil described in § 5.1.1. By tracking all these fusion terms, we see that they correspond to adding $-\psi_{\underline{z}^{-}}$, for $\underline{z}^{-}$ given by $z_{\alpha\beta}^{-}=\frac{1}{2}$, $\alpha<\beta$. It follows from the proof of Proposition 5.6 that these are precisely the terms that are killed to go from {,}$\{-,-\}_{0}$ (the local expression of the bivector $P$) to {$\{-,-\}_{-}$ (the Poisson structure from [5]). Forgetting the fusion of these $\mathbb{C}^{\times}$ actions can be reinterpreted as follows. We consider $\overline{Q}_{d}^{\infty_{d}}$, where $Q_{d}^{\infty_{d}}$ consists of the vertex set $\{0,\infty_{1},\ldots,\infty_{d}\}$ and the arrows $x:0\to 0$, $v_{\alpha}:\infty_{\alpha}\to 0$ for $1\leq\alpha\leq d$. We then perform the reduction of $M_{\overline{Q}_{d}^{\infty_{d}}}^{(0,1,\ldots,1)}(n,1,\ldots,1)$ at $q\operatorname{Id}_{n}\in\operatorname{GL}_{n}(\mathbb{C})$, i.e. we consider $\infty_{1},\ldots,\infty_{d}$ as ‘framing vertices’ with respect to which we are not reducing [^14] We use the standard trick that a (Hamiltonian) quasi-Poisson $(G\times H)$-variety where $H$ is abelian can be viewed as a (Hamiltonian) quasi-Poisson $G$-variety. Here $H=(\mathbb{C}^{\times})^{d}$ and $G=\operatorname{GL}_{n}(\mathbb{C})$.. In this way, Van den Bergh’s quasi-Poisson structure associated with $Q_{d}^{\infty_{d}}$ yields the minus Poisson structure of [5] in local coordinates.
@@ 5.2.3. Hamiltonian formulation of the spin RS system
[…]
@@
For any $\underline{z}\in\mathbb{C}^{d(d-1)/2}$, let {,}:={,}$\{-,-\}:=\{-,-\}_{0}+\psi^{\mathrm{loc}}_{\underline{z}}$. Then the equations (5.26a) hold for $h:=2(q^{-1}-1)\operatorname{tr}(Z)=2\sum_{i}f_{ii}$. In particular, any non-degenerate Poisson bracket from the pencil on $\mathfrak{h}_{\mathrm{reg}}/S_{n}$ gives the equations (5.26a), up to rescaling of $h$.
@@
Note that $\psi^{\mathrm{loc}}_{\underline{z}}(h,-)=0$ because $h$ is a linear combination of Casimirs by Lemma 5.3. Thus $\{h,-\}=\{h,-\}_{0}$, and (5.26a) holds for {,}$\{-,-\}_{0}$ by [12].
For the second part, recall that a non-degenerate Poisson bracket from the pencil is of the form {,}$\{-,-\}=z_{0}\{-,-\}_{0}+\psi^{\mathrm{loc}}_{\underline{z}}$ with $z_{0}\in\mathbb{C}^{\times}$. In that case we get (5.26a) for $\frac{d}{dt}=\{z_{0}^{-1}h,-\}$. ∎

### v3

@@ 4. Noncommutative origin of the pencil
As a consequence of Theorem 4.5, we can repeat Remark 4.2 and see the family of Poisson brackets $(z_{0}\{-,-\}_{\mathtt{VdB}}+\{-,-\}_{\underline{z}})$ on $\operatorname{Sym}(H_{0}(A_{Q}))$ as the universal Poisson pencil that induces the one on the varieties $M_{\overline{Q}}^{\gamma}/\!/\operatorname{GL}_{\mathbf{n}}$ defined by taking invariants in Theorem 3.14. In fact, any computation can be performed using the operations $(\left\{\!\!\left\{-,-\right\}\!\!\right\}_{o},\left\{\!\!\left\{-,-\right\}\!\!\right\}_{r})$ on $A_{Q}$.
@@
An analogue of Theorem 4.5 in the case of character varieties (cf. § 3.3.2) can be obtained by replacing Van den Bergh’s outer double (quasi-Poisson) bracket on $A_{Q}$ by the one introduced by Massuyeau-Turaev in [31] on $\mathbb{C}\langle a_{1}^{\pm 1},(a_{1}^{\ast})^{\pm 1},\ldots,a_{g}^{\pm 1},(a_{g}^{\ast})^{\pm 1},z_{1}^{\pm 1},\ldots,z_{r}^{\pm 1}\rangle$. The case of quiver varieties (cf. § 3.3.4) is obtained similarly by using Van den Bergh’s outer double Poisson bracket on $\mathbb{C}\overline{Q}$ given in [39, §6.3-6.4]. Alternatively note that the outer double bracket on $\mathbb{C}\overline{Q}$ in the Poisson case reduces to the first constant term appearing in (4.2), thus we can formally derive this case from Theorem 4.5 by setting $\gamma_{a}=\tilde{\gamma}\in\mathbb{C}^{\times}$ for each $a\in\overline{Q}$ and considering the family $(z_{0}\tilde{\gamma}^{-1}\{-,-\}_{\mathtt{VdB}}+\{-,-\}_{\underline{z}})$ in the limit $\tilde{\gamma}\to\infty$. For generalized MQV (cf. § 3.3.3), it is conjectured in [19] that their Poisson structure is induced by an outer double bracket; as this problem is still open, exhibiting a universal Poisson pencil is currently beyond reach.
@@ 5.1. General construction
$M_{\overline{Q}_{d}}^{\gamma}$= $\{(X,Z,V_{\alpha},W_{\alpha})\mid\det(X),\det(Z),\det(\operatorname{Id}_{n}+W_{\alpha}V_{\alpha})\neq 0,1\leq\alpha\leq d\}$ (5.1)
@@
$M_{\overline{Q}_{d}}^{\gamma}=$ $\{(X,Z,A_{\alpha},B_{\alpha})\mid\det(X),\det(Z),\det(\mathcal{Z}_{\alpha})\neq 0,1\leq\alpha\leq d\}$ (5.6)
@@ 5.1.1. The pencil
Due to Theorem 3.14, the bivector $P+\psi_{\underline{z}}$ is always non-degenerate quasi-Poisson on $M_{\overline{Q}_{d}}^{\gamma}$ for the moment map (5.3) (or equivalently (5.7)). In particular, this yields a quasi-Poisson pencil centered at $P$ of order d(d1)/2d(d-1)/2. After performing reduction, we obtain a Poisson pencil on the spin RS phase space $\mathcal{C}_{n,d,q}^{\times}$ (5.8).
@@ 5.2. Local parametrization
We continue with the notations of the previous subsection. Following [12, §4.1], consider the affine space $(\mathbb{C}^{\times})^{n}\times\mathbb{C}^{2nd}$ with coordinates $(x_{i},a_{i}^{\alpha},b_{i}^{\alpha})_{1\leq i\leq n}^{1\leq\alpha\leq d}$, and define the following subspace of dimension $2nd$:
@@
is well-defined and injective. Furthermore, $\xi$ is a morphism of Poisson varieties if $\mathcal{C}_{n,d,q}^{\times}$ is endowed with the Poisson bracket defined by the Poisson bivector $P$ (3.16) and $\mathfrak{h}_{\mathrm{reg}}/S_{n}$ is endowed with the Poisson bracket $\{-,-\}_{0}$ uniquely determined by
@@
$f_{k}:=\operatorname{tr}(X^{k}),\quad g_{\alpha\beta;k}:=\operatorname{tr}(A_{\alpha}B_{\beta}X^{k}),\quad k\geq 0,\,\,1\leq\alpha,\beta\leq d,$ (5.14)
@@ 5.2.1. Local expression of $\psi_{\underline{z}}$
By reduction of the quasi-Poisson pencil given in § 5.1.1, any bivector $z_{0}P+\psi_{\underline{z}}$ on $\mathcal{C}_{n,d,q}^{\times}$ with $z_{0}\in\mathbb{C}$ is Poisson, and it is non-degenerate whenever $z_{0}\neq 0$, cf. Corollary 3.10. Restricting the Poisson pencil to $\mathcal{C}^{\circ}$, the isomorphism $\xi:\mathfrak{h}_{\mathrm{reg}}/S_{n}\to\mathcal{C}^{\circ}$ that intertwines the Poisson structures $\{-,-\}_{0}$ with $P$ (by Proposition 5.1) and $\psi_{\underline{z}}^{\mathrm{loc}}$ with $\psi_{\underline{z}}$ (by Lemma 5.2) is therefore intertwining the structure $z_{0}\{-,-\}_{0}+\psi_{\underline{z}}^{\mathrm{loc}}$ with $z_{0}P+\psi_{\underline{z}}$, making the former a Poisson bracket for any possible parameter. These observations can be summarized in the following form.
@@
There is a Poisson pencil on $\mathfrak{h}_{\mathrm{reg}}/S_{n}$ spanned by the Poisson brackets $\{-,-\}_{0}$ and $\psi_{\underline{z}}^{\mathrm{loc}}$, with $\underline{z}=(z_{\alpha\beta})_{1\leq\alpha<\beta\leq d}$ defined as in § 5.1.1. Any linear combination $z_{0}\{-,-\}_{0}+\psi_{\underline{z}}^{\mathrm{loc}}$ with $z_{0}\in\mathbb{C}$ is therefore a Poisson bracket, which is non-degenerate when $z_{0}\neq 0$. This pencil (and therefore the pencil spanned by $P$ and $\psi_{\underline{z}}$ on $\mathcal{C}_{n,d,q}^{\times}$) has order $r\geq\frac{(d-1)(d-2)}{2}+1$.
@@
We only need to compute the bound on the order of the pencil. Since each $\psi_{\underline{z}}^{\mathrm{loc}}$ is degenerate while $\{-,-\}_{0}$ is not, the order equals $\mathrm{ord}(\psi_{\underline{z}}^{\mathrm{loc}})+1$.
@@
Remark $5.$
We expect that the inequality for the order given in Corollary 5.4 is an equality. This is the case for $d=2$, as the order is precisely 11. To see this, note that $\psi_{\underline{z}}^{\mathrm{loc}}\equiv 0$ identically for any $z_{12}\in\mathbb{C}$ because $G_{\underline{z}}(i,j;\alpha,\beta)$ (5.17) is always vanishing on $\mathfrak{h}_{\mathrm{reg}}$.
@@ 5.2.2. Comparison with the structure of Arutyunov-Olivucci
$\pm\frac{1}{2}\sum_{\gamma=1}^{d}o(\alpha,\gamma)b_{j}^{\beta}a_{j}^{\alpha}a_{i}^{\gamma}\pm\frac{1}{2}\left(\sum_{\gamma,\epsilon}o(\epsilon,\gamma)a_{i}^{\epsilon}a_{j}^{\gamma}\right)a_{i}^{\alpha}b_{j}^{\beta}+\frac{1}{2}a_{i}^{\alpha}a_{i}^{\beta}b_{j}^{\beta}-\frac{1}{2}\delta_{\alpha\beta}a_{i}^{\alpha}b_{j}^{\beta}\,$, (5.24c)
@@
The minus Poisson bracket $\{-,-\}_{-}$ constructed by Arutyunov and Olivucci [5] and the Poisson bracket $\{-,-\}_{0}$ constructed by Chalykh and Fairon [12] belong to the pencil of compatible Poisson brackets from Corollary 5.4.
@@
By definition, $\{-,-\}_{0}$ is certainly in the pencil. Take the sequence of parameters $\underline{z}^{-}$ given by $z_{\alpha\beta}^{-}=\frac{1}{2}$, $\alpha<\beta$. The corresponding skewsymmetric matrix has entries $\frac{1}{2}o(\alpha,\beta)$, and the Poisson structure $\psi_{\underline{z}^{-}}^{\mathrm{loc}}$ (5.16) is given in terms of
@@
It is then an exercise to check that $\{-,-\}_{-}=\{-,-\}_{0}+\psi_{\underline{z}^{-}}^{\mathrm{loc}}$ using (5.13) and (5.24), remembering that $\sum_{\gamma}a_{j}^{\gamma}=1$ for each $1\leq j\leq n$. ∎
@@
Finally, we can use Proposition 5.6 to provide another quiver interpretation of the minus Poisson structure of Arutyunov and Olivucci [5]. To do so, we need to unwind the construction of the master phase space $M_{\overline{Q}_{d}}^{\gamma}$ considered in § 5.1. At a purely geometric level, the quasi-Poisson structure is obtained by considering the internally fused double $\operatorname{GL}_{n}\times\operatorname{GL}_{n}=\{(X,Z)\}$ as a $\operatorname{GL}_{n}$-variety, with $d$ copies of the space $(T^{\ast}\mathbb{C}^{n})^{\circ}:=\{(W,V)\in T^{\ast}\mathbb{C}^{n}\mid 1+VW\neq 0\}$ seen as a $(\operatorname{GL}_{n}\times\mathbb{C}^{\times})$-variety, and we perform fusion of these different actions in a specific order. Observe from Remark 3.7 that the fusion terms added by the $\mathbb{C}^{\times}$ actions will, in fact, appear as bivectors $\psi_{\underline{z}}$ in the pencil described in § 5.1.1. By tracking all these fusion terms, we see that they correspond to adding $\psi_{\underline{z}^{-}}$, for $\underline{z}^{-}$ given by $z_{\alpha\beta}^{-}=\frac{1}{2}$, $\alpha<\beta$. It follows from the proof of Proposition 5.6 that these are precisely the terms that are killed to go from $\{-,-\}_{0}$ (the local expression of the bivector $P$) to $\{-,-\}_{-}$ (the Poisson structure from [5]). Forgetting the fusion of these $\mathbb{C}^{\times}$ actions can be reinterpreted as follows. We consider $\overline{Q}_{d}^{\infty_{d}}$, where $Q_{d}^{\infty_{d}}$ consists of the vertex set $\{0,\infty_{1},\ldots,\infty_{d}\}$ and the arrows $x:0\to 0$, $v_{\alpha}:\infty_{\alpha}\to 0$ for $1\leq\alpha\leq d$. We then perform the reduction of $M_{\overline{Q}_{d}^{\infty_{d}}}^{(0,1,\ldots,1)}(n,1,\ldots,1)$ at $q\operatorname{Id}_{n}\in\operatorname{GL}_{n}(\mathbb{C})$, i.e. we consider $\infty_{1},\ldots,\infty_{d}$ as ‘framing vertices’ with respect to which we are not reducing [^14] We use the standard trick that a (Hamiltonian) quasi-Poisson $(G\times H)$-variety where $H$ is abelian can be viewed as a (Hamiltonian) quasi-Poisson $G$-variety. Here $H=(\mathbb{C}^{\times})^{d}$ and $G=\operatorname{GL}_{n}(\mathbb{C})$.. In this way, Van den Bergh’s quasi-Poisson structure associated with $Q_{d}^{\infty_{d}}$ yields the minus Poisson structure of [5] in local coordinates.
@@ 5.2.3. Hamiltonian formulation of the spin RS system
It is proved in [5, 12] that, with $h:=2(q^{-1}-1)\operatorname{tr}(Z)$, the equations of motion for the derivation $\frac{d}{dt}:=\{h,-\}$ (where $\{-,-\}$ denotes the Poisson bracket from the corresponding reference) satisfy locally
@@
For any $\underline{z}\in\mathbb{C}^{d(d-1)/2}$, let $\{-,-\}:=\{-,-\}_{0}+\psi^{\mathrm{loc}}_{\underline{z}}$. Then the equations (5.26a) hold for $h:=2(q^{-1}-1)\operatorname{tr}(Z)=2\sum_{i}f_{ii}$. In particular, any non-degenerate Poisson bracket from the pencil on $\mathfrak{h}_{\mathrm{reg}}/S_{n}$ gives the equations (5.26a), up to rescaling of $h$.
@@
Note that $\psi^{\mathrm{loc}}_{\underline{z}}(h,-)=0$ because $h$ is a linear combination of Casimirs by Lemma 5.3. Thus $\{h,-\}=\{h,-\}_{0}$, and (5.26a) holds for $\{-,-\}_{0}$ by [12].
For the second part, recall that a non-degenerate Poisson bracket from the pencil is of the form $\{-,-\}=z_{0}\{-,-\}_{0}+\psi^{\mathrm{loc}}_{\underline{z}}$ with $z_{0}\in\mathbb{C}^{\times}$. In that case we get (5.26a) for $\frac{d}{dt}=\{z_{0}^{-1}h,-\}$. ∎